$2500 is invested at 6% per year compounded annually. How long will it take this investment to grow to $10,000?
Answer by addingup(3677)

2500(1+0.06)^t = 10000
1.06^t = 4
tlog(1.06) = log(4)
t = log(4)/log(1.06)
t = 23.8 years
------------------------------------
Check:
2500(1+.06)^23.8 = 2500(1.06)^23.8 = 2500*4 = 10000 Correct